### v1

@@ 7 Acknowledgements
We thank Davor Tolj, Tiffany Soetojo and Rachel Michel for helpful discussions. This work made use of the bulk crystal growth facility of the NSF’s Platform for the Accelerated Realization, Analysis, and Discovery of Interface Materials (PARADIM), which is supported by the National Science Foundation under Cooperative Agreement No. DMR-2039380. This work also utilized the JHU Raman Scattering Users Center. Work at UC Santa Barbara used facilities supported via the UC Santa Barbara NSF Quantum Foundry funded via the Q-AMASE-i program under award number DMR-1906325. This work was also supported by the BONASPES (ANR-19-CE30-0007) and HETERO-BNC (ANR-20-CE09-0014-02) projects. The MPMS was funded by the National Science Foundation, Division of Materials Research, Major Research Instrumentation Program, under award No. 1828490. GY, CN, and RH acknowledge support from NSF Grants No. DMR-2300640 and No. DMR-2104036. SDW and RG acknowledge support from the Air Force Office of Scientific Research under award No. FA9550-23-1-0042. JHE acknowledges support from the Office of Naval Research award No. N00014-22-1-2582. EZ recognizes financial support from JHU through the Sweeney Family Postdoctoral Fellowship.
8 Data availability statement
All data collected in PARADIM facilities to carry out this work will be available upon publication at https://doi.org/ xxxx-xxxx. Other data is available upon reasonable request to the authors.

### v2

@@ 7 Acknowledgements
We thank Davor Tolj, Tiffany Soetojo and Rachel Michel for helpful discussions. This work made use of the bulk crystal growth facility of the National Science Foundation’s Platform for the Accelerated Realization, Analysis, and Discovery of Interface Materials (PARADIM), which is supported by the National Science Foundation under Cooperative Agreement No. DMR-2039380. This work also utilized the JHU Raman Scattering Users Center. Work at UC Santa Barbara used facilities supported via the UC Santa Barbara National Science Foundation Quantum Foundry funded via the Q-AMASE-i program under award number DMR-1906325. This work was also supported by the BONASPES (ANR-19-CE30-0007) and HETERO-BNC (ANR-20-CE09-0014-02) projects. The MPMS was funded by the National Science Foundation, Division of Materials Research, Major Research Instrumentation Program, under award No. 1828490. GY, CN, and RH acknowledge support from National Science Foundation Grants No. DMR-2300640 and No. DMR-2104036. SDW and RG acknowledge support from the Air Force Office of Scientific Research under award No. FA9550-23-1-0042. JHE acknowledges support from the Office of Naval Research award No. N00014-22-1-2582.
8 Data availability